OpenStudy (anonymous):

28^2 in standard notation? 17^3 in standard notation? 8^3 in standard notation? 6^4 in standard notation? 5^4 in standard notation?

OpenStudy (anonymous):

Bipasha, Use your calculator The exponent means repeated how many times: 28 * 28 = 784 17 * 17 * 17 = 4913 8 * 8 * 8 = 6* 6 * 6 * 6 = ...